OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10 I got x=-3 and x=1 Please check my answer!

Nnesha (nnesha):

nah what was your first step ?

OpenStudy (anonymous):

dividing both sides by -5

Nnesha (nnesha):

yep then what did you get at right side ?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

i subtracted 1 from both sides and got x=-3

OpenStudy (johnweldon1993):

So then you have \[\large |x + 1| = -2\] So just looking at that equation...think...does an absolute value ever equal a negative number?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

So it's no solution?

OpenStudy (johnweldon1993):

So there are actually no solutions to this expression

OpenStudy (anonymous):

Thank you!

Nnesha (nnesha):

yep that's right absolute is always positive \[\left| x \right| \ge 0\]

Nnesha (nnesha):

value*

OpenStudy (anonymous):

Thank you, @Nnesha and @johnweldon1993 ! I fanned both of you.